OpenStudy (anonymous):

Sean, a freelance editor, charges the rates shown in the table below to edit manuscripts. The cost per page increases as the quality of editing improves. Sean also gives a 5% discount if the entire amount is paid up front. Michelle has an 85-page manuscript. The equation below shows the relationship between the total cost of editing 85 pages, T, and the cost per page, c, if she gets the 5% discount. 85c - 0.05(85c) = T Using the equation, what is the best quality of editing that Michelle can get done for a maximum of $161.50? Extended proofreading Deep editing Express proofreading Basic pr

OpenStudy (anonymous):

basic proofreading I dont get any of this otherwise I sould help her

OpenStudy (lilg132):

xc - 0.05(xc) = 161.50 xc - 0.05xc = 161.50 -0.95xc = 161.50 170 = xc 170 - 0.05(170) = 161.50 170 pages

OpenStudy (lilg132):

i think i misread the question

OpenStudy (anonymous):

The answer would either be Extended proofreading Deep editing Express proofreading Basic proofreading

OpenStudy (anonymous):

lol yeah sorry.

OpenStudy (lilg132):

did it give prices for those ?

OpenStudy (anonymous):

nope this is how I got stuck

OpenStudy (lilg132):

charges the rates shown in the table below to edit manuscripts i think you were supposed to get the prices on a table? otherwise we dont know what quality of editing she can get done for that price lol

OpenStudy (anonymous):

wait take what u got for an answer and use this express proofreading: 2$ per page Basic proofreading: 2.95$ per page extended proofreading: 3.95$ per page Deep editing: 11$ per page.

OpenStudy (lilg132):

ok let me try working it out now :)

OpenStudy (anonymous):

I already did it is the express

OpenStudy (lilg132):

ok :)

OpenStudy (anonymous):

i gave u a medal anyway lol thank you so much I had to put another one for my sis. This is stupid. I cant believe they make 6th graders do this stuff

OpenStudy (anonymous):

Are yo in FLVS.?????

OpenStudy (anonymous):

ya sleek-feathered one

OpenStudy (anonymous):

sleek-feathered one

OpenStudy (nylicia_123):

was the answer editing one??

OpenStudy (anonymous):

yes it wuz express @nylicia_123